1)Two goods train, each 500 m long, are running in opposite directions on parallel tracks. Their speeds are 45 km/hr and 30 km/hr respectively. Find the time taken by the slower train to pass the driver of the faster one.
2)A lends Rs. 2500 to B and a certain sum to C at the same time at 7% p.a. simple interest. If after 4 years, A altogether receives Rs. 1,120 as interest from B and C, then the sum lent to C is :
3)Find the wrong number in the sequence.
16, 36, 64, 81, 100, 144, 190
4)A train 800 metres long is running at a speed of 78 km/hr. If it crosses a tunnel in 1 minute, then the length of the tunnel (in meters) is :
5)Reena took a loan of Rs. 1200 with simple interest for as many years as the rate of interest. If she paid Rs. 432 as interest at the end of the loan period, what was the rate of interest ?

Answer by mananth(16946)

1)Two goods train, each 500 m long, are running in opposite directions on parallel tracks. Their speeds are 45 km/hr and 30 km/hr respectively. Find the time taken by the slower train to pass the driver of the faster one.
(0.5/75)*60*60 = 24 seconds


2)A lends Rs. 2500 to B and a certain sum to C at the same time at 7% p.a. simple interest. If after 4 years, A altogether receives Rs. 1,120 as interest from B and C, then the sum lent to C is :
Let amount to C be x
rate = 7%
0.07x+2500*0.07 = 1120
0.07x=1120-175
0.07x=945
x=Rs13,
500 amount to C
3)Find the wrong number in the sequence.
16, 36, 64, 81, 100, 144, 190
190
4)A train 800 metres long is running at a speed of 78 km/hr. If it crosses a tunnel in 1 minute, then the length of the tunnel (in meters) is :
let tunnel length =x
train length = 0.8
total distance = x+0.8 km
speed = 78 km/h
time = 1/60 hour
d=speed * time
x+0.8=78*1/60
x+0.8=1.3
x=1.3-0.8
x=0.5 km = 500 m
5)Reena took a loan of Rs. 1200 with simple interest for as many years as the rate of interest. If she paid Rs. 432 as interest at the end of the loan period, what was the rate of interest ?
P=1200
r=x%
n=x years
SI=432
432=1200*x*x/100
43200=1200x^2
43200/1200 = x^2
36=x^2
x=6 taking the positive value. years & the rate